A 3.4 μC charge is placed at the origin of coordinates, and a -2.6 μC charge is placed to the x-axis at 4.0 cm.

Part A.

Find the location of the place(s) along the x axis where the electric field due to these two charges is zero.

Part B

Find the location of the place(s) along the x axis where the electric potential due to these two charges is zero.

Please show the work. Thank you. Constants A 3.4 uC charge is placed at the origin of coordinates, and a -2.6 C charge is placed to the I axis at 4.0 cm Part
Answer #1

Xo X-4 cm q=3-40c ممے q= -2.6 kic (4772 쁨 Hr 3.4 2-6 8= 27.86 x = 32cm Up = 0; Verzo kaut keho 7 (8+4) 8 2.6 3.4 34 8=2.430w